OpenStudy (loser66):

Prove that in an ordered field , if \(\sqrt 2\) is a positive number whose square is 2, then \(\sqrt 2<3/2 \) Please, help

OpenStudy (loser66):

@SithsAndGiggles

OpenStudy (loser66):

Suppose \(\sqrt 2>3/2\) then \(\sqrt 2−3/2>0\) square a positive number gives us a positive number also. So that \(\sqrt2−3/2)^2>0\) it gives me \(2−3\sqrt 2+9/4>0\) that means \(17/4>3\sqrt 2\) or \(17/12>\sqrt 2\) since both sides are >0, we square it , the sign doesn't change and I get 2.0006 > 2 I go nowhere. ;(

OpenStudy (loser66):

However, I don't see my mistake, please, point it out.

OpenStudy (loser66):

Oh, I got it, friend. :) Thanks for being here